OpenStudy (anonymous):

A reaction that is first order with respect to reactant A has a rate constant 6 min^-1. If we start with [A]=50mol/L, when would [A] reach the value 0.05mol/L.

OpenStudy (frostbite):

Just insert into the equation. You have all the information: \[[A]=[A _{0}]e ^{-kt}\] Isolate t: \[t=\frac{ -\ln(\frac{ [A] }{ [A _{0}] }) }{ k }\]

OpenStudy (frostbite):

Do a check for the units your self ;)

OpenStudy (anonymous):

plz solve using \[t=\frac{ 2.303 }{ k }\log_{10}\frac{ A_{0} }{ A } \]

OpenStudy (frostbite):

Also a option... just insert into the equation.

OpenStudy (anonymous):

ya so what is A and A0 here.......i am not getting the answer i request you to solve the question step by step and show the solution. I have tried doing the sum 5 times can u plz help me.......?

OpenStudy (frostbite):

Okay. I help you the 0 in A0 is a reference to the following A at t=0... so A0 = 50mol/L.... A=0.05mol/L. k is rate constant and so on... in other words: \[t=\frac{ -\ln(\frac{ [A] }{ [A _{0}] }) }{ k }=\frac{ -\ln(\frac{ [0.05] }{ 50 }) }{ 6 }\ I've been lazy and not put the units on the calculations... you should do so.

OpenStudy (frostbite):

\[t=\frac{ -\ln(\frac{ [A] }{ [A _{0}] }) }{ k }=\frac{ -\ln(\frac{ [0.05] }{ 50 }) }{ 6 }\]

OpenStudy (anonymous):

I am Sorry but after substituting the values we get 0.5min as the answer but the solution says it is 0.77min. Please check......

OpenStudy (frostbite):

I don't see how you get 0.5... nor 0.77 I want it to be 1 min and 15 sec. \[t=\frac{ -\ln(\frac{ [0.05] }{ 50 }) }{ 6 }=\frac{ -\ln (0.001) }{ 6 } \approx \frac{ -(-6.90776) }{ 6 } \approx 1.15129\]

OpenStudy (frostbite):

sorry not 1 min and 15 sec but 1.15 min.

OpenStudy (chmvijay):

yaaa its 1.15 :) i too checked thrice

OpenStudy (chmvijay):

hey please check the question is it 0.05 or 0.5 mol/L if we take 0.5 ur desired 0.77 min answer we get:)

OpenStudy (chmvijay):

@leena1996

OpenStudy (anonymous):

Alright...thanks..........@chmvijay @Frostbite

OpenStudy (frostbite):

Nice noticed chmvijay.

OpenStudy (chmvijay):

humnn thank you @Frostbite and @leena1996 :)

OpenStudy (frostbite):

Well back to electrochemistry for me... I get too easy distracted.

OpenStudy (chmvijay):

:)

OpenStudy (anonymous):

@Frostbite Oops sorry.....get back to Work....!

OpenStudy (dean.shyy):

This may help: http://is.gd/Ytba92 & http://is.gd/auHUsw